# Semiformal method I. Pólya’s theorem for the complete graph $K_{\mathbb{N}}$ with complex edge weights

### v2

@@ 1 Introduction
Pólya’s theorem is almost always cited in the special case with $\overline{v}=\overline{0}$, although G. Pólya proved it for any vertex $\overline{v}$ (see [15, Section 1] for a translated quote from [23]). A selection of articles and books discussing Pólya’s theorem is [3, 4, 6, 9, 11, 13, 17, 18, 19, 20, 24, 25, 28, 29].
@@
be the edges of the countable complete graph $K_{\mathbb{N}}=\langle\mathbb{N},\mathbb{N}_{2}\rangle$ on the vertex set $\mathbb{N}$. For a finite set $A$ we denote by $|A|$ ($\in\mathbb{N}_{0}=\{0,1.\dots\}$) its cardinality. Let
@@
We treat this series both in conditional and absolute convergence. Our formulas appear in fourteen theorems: Theorems 3.23.7 in Section 3 and Theorems 4.24.9 in Section 4. Pólya’s theorem is subsumed in five of them: Theorems 3.5 and 3.7 extend cases $d\geq 3$ and $d\leq 2$, respectively, when $\overline{v}=\overline{0}$, and Theorems 4.6 (4.7) and 4.9 cases $d\geq 3$ and $d\leq 2$, respectively, when $\overline{v}\neq\overline{0}$.
Why fourteen generalizations of one theorem? For three reasons. First, in Pólya’s theorem weights are nonnegative real numbers and convergence of infinite series is automatically absolute. For complex weights we consider both absolute, and conditional convergence. We also consider, for nonnegative real weights, series with sum $\infty$. Second, Pólya’s theorem lives in the probabilistic setup where weights of edges incident to a fixed vertex sum up to $1$. We consider two cases—when this condition is kept (convex weights) and when it is dropped. Third, already Pólya’s theorem is in reality two theorems because the cases $\overline{v}=\overline{0}$ and $\overline{v}\neq\overline{0}$ differ in the complexity of proofs. In our generalization we correspondingly distinguish cases $v=1$ and $v\neq 1$. Now $3\cdot 2\cdot 2=12$ and two more generalizations arise by variations of assumptions.
@@
the image of $C$ by $f$. Let $d\in\mathbb{N}$ and $\overline{v}\in\mathbb{Z}^{d}$. We take any bijection
$f\colon\mathbb{Z}^{d}\to\mathbb{N}$
such that $f(\overline{0})=1$, and set $v=f(\overline{v})$. The grid graph on $\mathbb{Z}^{d}$, i.e. the graph $\langle\mathbb{Z}^{d},E_{d}\rangle$ with the edges
@@
We structure our article as follows. In Section 2 we review infinite series with complex summands and generating functions. Sections 3 and 4 contain our main results, the mentioned fourteen generalizations of Pólya’s theorem. Section 5 contains concluding remarks. We continue our investigation in [16] where we consider edge weights
$h\colon\mathbb{N}_{2}\to\mathbb{C}[[x_{1},\,\dots,\,x_{k}]]\,.$
The domain of formal power series $\mathbb{C}[[x_{1},\dots,x_{k}]]$ is endowed with the standard non-Archimedean norm.
In Sections 3 and 4 we use what we term the semiformal method in (enumerative) combinatorics. It extends the symbolic method in combinatorics [2, 8, 10, 12, 21, 22, 26] from finite to countable sets. In our situation such extension is natural and inevitable, coefficients in generating functions $A_{h,v}(x)$ arise as sums of infinite series $\sum_{w\in W(v,n)}h(w)$. We develop the semiformal method more in [14]. Why “semiformal”? Generating “functions” in the symbolic method are formal power series with coefficients determined by finite operations. Our extension allows operations with coefficients based on limit transitions.
2 Series and generating functions
We review series and generating functions. A series is a map
$h\colon X\to\mathbb{C}$
defined on an at most countable set $X$. We write it as $\sum_{x\in X}h(x)$. We say that $\sum_{x\in X}h(x)$ absolutely converges if two equivalent conditions hold.
@@ 2 Series and generating functions
such that $n\in\mathbb{N}_{0}$ and $v_{i-1}\neq v_{i}$ for every $i\in[n]$ ($\{1,2,\dots,n\}$, $0]=\emptyset$). The length $n$ of $w$ is denoted by $|w|$ ($\in\mathbb{N}_{0}$). We extend $h$ to walks. For $n>0$ we set
@@
and for $n=0$ we define $h(w)=1$. Recall that $W(n)$ is the set of walks $w$ in $K_{\mathbb{N}}$ such that $v_{0}=1$ and $|w|=n$.
@@
Let $S=\sum_{x\in X}h(x)$ be a series. If $Y\subset X$, we say that $\sum_{x\in Y}h(x)$ is a subseries of $S$. Any subseries of an absolutely convergent series absolutely converges, but we cannot compare sizes of sums
@@
Except when $h\colon X\to\mathbb{R}_{\geq 0}$. If the series absolutely converges and $Y\subset X$, then the sums satisfy
@@
If $S=\sum_{x\in X}h(x)$ absolutely converges and has sum $s$, and $P$ is a partition of $X$, then $S_{P}$ is correctly defined, absolutely converges and has the same sum $s$.
Proof. For every $Z\in P$ the series $S_{Z}=\sum_{x\in Z}h(x)$ is a subseries of $S$ and absolutely converges. Thus the series $S_{P}$ is correctly defined and we show that it absolutely converges. Let
@@
Let $\{Z_{1},Z_{2},\dots,Z_{n}\}\subset P$ be a finite set. We use Lemma 2.3 and take for each $i\in[n]$ the finite set
@@
Proof. We show that $S_{1}\cdot S_{2}$ absolutely converges. Let $c\geq 0$ be such that for any finite sets $Y_{i}$ with $Y_{i}\subset X_{i}$ and $i=1,2$,
$\sum_{x\in Y_{i}}|h_{i}(x)|\leq c\,.$}
Let $A\subset X_{1}\times X_{2}$ be a finite set. We take any finite sets $Y_{i}\subset X_{i}$ such that $A\subset Y_{1}\times Y_{2}$ and get the desired bound:
@@
where
[…]
Hence $|s-s_{1}s_{2}|\leq\frac{\varepsilon}{4}+|s_{1}t_{2}|+|s_{2}t_{1}|+|t_{1}t_{2}|\leq\frac{\varepsilon}{4}+\frac{\varepsilon}{4}+\frac{\varepsilon}{4}+\frac{\varepsilon}{4}=\varepsilon$. $\Box$
Let $S_{i}=\sum_{x\in X_{i}}h_{i}(x)$, $i=1,2,\dots,k$, be series. We define more generally their product as the series
@@
If the series $X_{1}$, $\dots$, $X_{k}$ absolutely converge and have sums $s_{1}$, $\dots$, $s_{k}$, respectively, then $\prod_{i=1}^{k}S_{i}$ absolutely converges and has the sum $s_{1}s_{2}\dots s_{k}$.
@@
If nonempty series $S_{1}$ and $S_{2}$ are not identically zero and if $S_{1}\cdot S_{2}$ absolutely converges and has the sum $s$, then $S_{1}$ and $S_{2}$ absolutely converge and their respective sums $s_{1}$ and $s_{2}$ satisfy $s_{1}s_{2}=s$
Proof. This follows from Proposition 2.5 if we show that $S_{1}$ and $S_{2}$ absolutely converge. Let $S_{i}=\sum_{x\in X_{i}}h_{i}(x)$. $i=1,2$, and let $c>0$ be such that for every finite set $X\subset X_{1}\times X_{2}$,
@@
We prove absolute convergence of $S_{1}$. We take any $y\in X_{2}$ with $h_{2}(y)\neq 0$. Then for every finite set $X\subset X_{1}$,
@@
Thus $S_{1}$ absolutely converges. For $S_{2}$ we proceed similarly. $\Box$
It is clear that the assumption on $S_{1}$ and $S_{2}$ cannot be omitted.
Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ and $n\in\mathbb{N}_{0}$. We obtain a sufficient condition for the lightness of $h$ in terms of the values $h(e)$. We denote by $V(h,n)$ the set of vertices in $K_{\mathbb{N}}$ that can be reached from $1$ by a walk of length at most $n$ that has edges with nonzero weights.
@@
Let $X\subset W(n)$ be finite. We may assume that walks in $X$ have only edges with nonzero weights. We decompose every walk $w\in X$ as $w=w^{\prime}\,\ell(w)$, where $w^{\prime}\in W(n-1)$ and $\ell(w)$ is the last vertex of $w$. Let $Y$ be the set of walks $w^{\prime}$. For $w^{\prime}\in Y$ let
@@
Using finite versions of Propositions 2.4 and 2.2 we get the bound
@@
Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$. We define
$V(h)=\bigcup_{n\geq 0}V(h,\,n)\,.$}
It is the set of vertices in $K_{\mathbb{N}}$ reachable from $1$ by a walk formed by edges with nonzero weights. We define the graph
@@
as the connected component of $1$ in the subgraph of $K_{\mathbb{N}}$ formed by edges with nonzero weight. It follows that if $G(h)$ has finite degrees then $h$ is slim.
@@
Proof. There is a walk $w_{0}=\langle u_{0},\dots,u_{n}\rangle$ in $G(h)$ such that $u_{0}=1$ and $u_{n}=u$. Let
$c=h(w_{0})=\prod_{i=1}^{n}h(\{u_{i-1},u_{i}\})\ \ (\neq 0)\,.$}
Let $W$ be the set of walks $\langle u_{0},\dots,u_{n},u_{n+1}\rangle$ in $K_{\mathbb{N}}$. Thus vertices $u_{0}$, $\dots$, $u_{n}$ are fixed and $u_{n+1}$ runs in $\mathbb{N}\setminus\{u_{n}\}=\mathbb{N}\setminus\{u\}$. The series $\sum_{w\in W}h(w)$ is a subseries of $\sum_{w\in W(n+1)}h(w)$ and therefore absolutely converges. By Proposition 2.2,
@@
We survey the results on generating functions, which we use in the next two sections. We work with univariate formal power series
@@
Neutral elements are $0=0+0x+0x^{2}+\dots$ and $1=1+0x+0x^{2}+\cdots$, and the units in $\mathbb{C}[[x]]$ are exactly the formal power series $\sum_{n\geq 0}a_{n}x^{n}$ with $a_{0}\neq 0$.
@@
If $U(1)_{2}$ exists, then $U(1)_{1}$ exists and $U(1)_{1}=U(1)_{2}$, but the opposite implication is, in general, not true. For example, for
@@
$U(1)_{1}$ exists and $U(1)_{1}=\log 2$, but $U(1)_{2}$ does not exist. But if
@@
If $U(x)=\sum_{n\geq 0}u_{n}x^{n}\in\mathbb{C}[[x]]$ and $U(1)_{1}$ exists, then for every number $x\in\mathbb{C}$ with $|x|<1$ the power series
$\sum_{n\geq 0}u_{n}x^{n}$}
absolutely converges.
@@
Quantities $U(1)_{1}$ are handled by means of the classical theorem of Abel, see [1] and [27, Chapter II.7]. We state it for the radius of convergence $R\geq 1$.
@@
Let $U(x),V_{1}(x),\dots,V_{n}(x)\in\mathbb{C}[[x]]$ for $n\in\mathbb{N}$. If $U(x)=\prod_{i=1}^{n}V_{i}(x)$ and $U(1)_{1}$, $V_{1}(1)_{1}$, $\dots$, $V_{n}(1)_{1}$ exist, then
@@
Proof. By Propositions 2.4, 2.6 and 2.13, and the assumptions,
$\text{$F_{U}(x)=\prod_{i=1}^{n}F_{V_{i}}(x)$ for every $x\in\mathbb{C}$ with $|x|<1$}\,.$}
Theorem 2.16 and properties of limits of functions give
@@
Proof. In $\mathbb{C}[[x]]$ we have the identity $U(x)V(x)=1$. By Propositions 2.2, 2.4, 2.5 and 2.13, $F_{U}(x)F_{V}(x)=1$ for $x\in[0,1)$. Thus $F_{V}(x)\neq 0$ and $F_{U}(x)=\frac{1}{F_{V}(x)}$ for every $x\in[0,\,1)$. Two applications of Theorem 2.17 and properties of limits of functions give
@@
Proof. In $\mathbb{C}[[x]]$ we have the identity $U(x)^{2}\cdot W(x)=W(x)-V(x)$. Let T(x)=W(x)V(x)T(x)=W(x)-V(x). By Proposition 2.18, the sum $T(1)_{1}$ exists and $T(1)_{1}=W(1)_{1}-V(1)_{1}$. By Proposition 2.19 we have $(U(1)_{1})^{2}\cdot W(1)_{1}=T(1)_{1}=W(1)_{1}-V(1)_{1}$. Division by $W(1)_{1}$ yields the stated formula. $\Box$
@@
$(U(1)_{1})^{2}$ $=$ ${\textstyle\lim_{x\to 1}F_{U}(x)^{2}=\lim_{x\to 1}\big(1-\frac{F_{V}(x)}{F_{W}(x)}\big)=1-\frac{\lim_{x\to 1^{-}}F_{V}(x)}{\lim_{x\to 1^{-}}F_{W}(x)}}$
$=$ ${\textstyle 1-\frac{a}{\lim_{n\to\infty}\sum_{j=0}^{n}w_{j}}=1-\frac{a}{+\infty}=1}\,.$
@@
We proceed to absolute convergence.
@@
It absolutely converges by Proposition 2.5 and has the sum $V(1)_{2}W(1)_{2}$. On the other hand, $S$ has the grouping $\sum_{n\geq 0}\sum_{j=0}^{n}v_{j}w_{n-j}$ (the inner $\sum$ means sum and the outer $\sum$ means series) and $\sum_{j=0}^{n}v_{j}w_{n-j}=u_{n}$. By Proposition 2.4, $S$ has the sum $u_{0}+u_{1}+\dots=U(1)_{2}$. Thus $U(1)_{2}=V(1)_{2}W(1)_{2}$.

### v3

@@ 1 Introduction
Pólya’s theorem is almost always cited in the special case with $\overline{v}=\overline{0}$, although he proved it for any vertex $\overline{v}$ (see [16, Section 1] for a translated quote from [24]). A selection of articles and books discussing Pólya’s theorem is [3, 4, 6, 9, 11, 13, 18, 19, 20, 21, 25, 26, 29, 30].
@@
be edges of the countable complete graph $K_{\mathbb{N}}=\langle\mathbb{N},\mathbb{N}_{2}\rangle$ on the vertex set $\mathbb{N}$. For a finite set $A$ we denote by $|A|$ ($\in\mathbb{N}_{0}=\{0,1.\dots\}$) its cardinality. Let
@@
The formulas appear in fourteen theorems: Theorems 3.23.7 in Section 3 and Theorems 4.24.9 in Section 4. Pólya’s theorem is subsumed in five of them. Theorems 3.5 and 3.7 extend cases $d\geq 3$ and $d\leq 2$, respectively, when $\overline{v}=\overline{0}$. Theorems 4.6 (4.7) and 4.9 extend cases $d\geq 3$ and $d\leq 2$, respectively, when $\overline{v}\neq\overline{0}$.
We have fourteen generalizations of one theorem for three reasons. First, in Pólya’s theorem weights are nonnegative real numbers and convergence of infinite series is automatically absolute. For complex weights we consider both kinds of convergence, absolute and conditional one. We also consider, for nonnegative real weights, series with sum $\infty$. Second, Pólya’s theorem lives in the probabilistic setup where weights of edges incident to a fixed vertex sum up to $1$. We consider the case when this condition is kept (convex weights) and the case when it is dropped. Third, already Pólya’s theorem is in reality two theorems because the cases $\overline{v}=\overline{0}$ and $\overline{v}\neq\overline{0}$ differ in complexity of proofs. In our generalization we distinguish cases $v=1$ and $v\neq 1$ as well. Now $3\cdot 2\cdot 2=12$; two more generalizations arise by variations of assumptions.
@@
the image of $C$ by $f$. Let $d\in\mathbb{N}$ and $\overline{v}\in\mathbb{Z}^{d}$. We take any bijection $f\colon\mathbb{Z}^{d}\to\mathbb{N}$ such that $f(\overline{0})=1$, and set $v=f(\overline{v})$. The grid graph on $\mathbb{Z}^{d}$, which is the graph $\langle\mathbb{Z}^{d},E_{d}\rangle$ with the edges
@@
We structure our article as follows. In Section 2 we review infinite series with complex summands and generating functions. Sections 3 and 4 contain our main results, the mentioned fourteen generalizations of Pólya’s theorem. Section 5 contains concluding remarks. We continue our investigation in [17] (see Section 5 for more details).
In Sections 3 and 4 we use what we term the semiformal method in (enumerative) combinatorics. It extends the symbolic method in combinatorics [2, 8, 10, 12, 22, 23, 27] from finite to countable sets. In our situation such extension is natural and inevitable, coefficients in generating functions $A_{h,v}(x)$ arise as sums of infinite series $\sum_{w\in W(v,n)}h(w)$. We develop the semiformal method more in [14]. Why “semiformal”? Generating “functions” in the symbolic method are formal power series with coefficients determined by finite operations. Our extension allows operations based on limit transitions. See [15] more applications of the semiformal method.
2 Series and generating functions in $\mathbb{C}$
We review series and generating functions in the domain $\mathbb{C}$. A series (in $\mathbb{C}$) is a map $h\colon X\to\mathbb{C}$, defined on an at most countable set $X$. We write it as $\sum_{x\in X}h(x)$. We say that $\sum_{x\in X}h(x)$ absolutely converges if two equivalent conditions hold.
@@ 2 Series and generating functions in $\mathbb{C}$
such that $n\in\mathbb{N}_{0}$ and $v_{i-1}\neq v_{i}$ for every $i\in[n]$ ($\{1,2,\dots,n\}$, $0]=\emptyset$). The length $n$ of $w$ is is $n$ ($\in\mathbb{N}_{0}$). We extend $h$ to walks. For $n>0$ we set
@@
and for $n=0$ we define $h(w)=1$. Recall that $W(n)$ is the set of walks $w$ in $K_{\mathbb{N}}$ with length $n$ such that $v_{0}=1$.
@@
Let $S=\sum_{x\in X}h(x)$ be a series. If $Y\subset X$, we say that $\sum_{x\in Y}h(x)$ is a subseries of $S$. Any subseries of an absolutely convergent series absolutely converges, but we cannot compare sizes of their sums
@@
Except when $h\colon X\to\mathbb{R}_{\geq 0}$: if $\sum_{x\in X}h(x)$ absolutely converges and $Y\subset X$, then the sums satisfy
@@
If $S=\sum_{x\in X}h(x)$ absolutely converges and has sum $s$, and if $P$ is a partition of $X$, then $S_{P}$ is correctly defined, absolutely converges and has the same sum $s$.
Proof. For every $Z\in P$ the series $S_{Z}=\sum_{x\in Z}h(x)$ is a subseries of $S$ and absolutely converges. Thus the series $S_{P}$ is correctly defined. We show that it absolutely converges. Let
@@
Let $\{Z_{1},Z_{2},\dots,Z_{n}\}\subset P$ be a finite set. We use Lemma 2.3 and take for $i\in[n]$ finite sets
@@
Proof. We show that $S_{1}\cdot S_{2}$ absolutely converges. Let $c\geq 0$ be such that for any finite sets $Y_{i}\subset X_{i}$, $i=1,2$, we have $\sum_{x\in Y_{i}}|h_{i}(x)|\leq c$. Let $A\subset X_{1}\times X_{2}$ be a finite set. We take any finite sets $Y_{i}\subset X_{i}$ such that $A\subset Y_{1}\times Y_{2}$ and get the desired bound:
@@
where $|t_{1}|\leq\frac{\varepsilon}{4(1+|s_{2}|)}$ and $|t_{2}|\leq\frac{\varepsilon}{4(1+|s_{1}|)}$. Hence $|s-s_{1}s_{2}|\leq\frac{\varepsilon}{4}+|s_{1}t_{2}|+|s_{2}t_{1}|+|t_{1}t_{2}|\leq\frac{\varepsilon}{4}+\frac{\varepsilon}{4}+\frac{\varepsilon}{4}+\frac{\varepsilon}{4}=\varepsilon$. $\Box$
Let $S_{i}=\sum_{x\in X_{i}}h_{i}(x)$, $i\in[k]$, be $k\in\mathbb{N}$ series. We define more generally their product as the series
@@
If the series $S_{1}$, $\dots$, $S_{k}$ absolutely converge and have sums $s_{1}$, $\dots$, $s_{k}$, respectively, then $\prod_{i=1}^{k}S_{i}$ absolutely converges and has sum $s_{1}s_{2}\dots s_{k}$.
@@
If nonempty series $S_{1}$ and $S_{2}$ are not identically zero and if their product $S_{1}\cdot S_{2}$ absolutely converges and has the sum $s$, then $S_{1}$ and $S_{2}$ absolutely converge and their respective sums $s_{1}$ and $s_{2}$ satisfy $s_{1}s_{2}=s$.
Proof. This follows from Proposition 2.5 if we show that $S_{1}$ and $S_{2}$ absolutely converge. Let $S_{i}=\sum_{x\in X_{i}}h_{i}(x)$, $i=1,2$, and let $c>0$ be such that for every finite set $X\subset X_{1}\times X_{2}$ we have
@@
We prove absolute convergence of $S_{1}$. We take any $y\in X_{2}$ with $h_{2}(y)\neq 0$. Then for every finite set $X\subset X_{1}$ we have
@@
Thus $S_{1}$ absolutely converges. For $S_{2}$ we argue similarly. $\Box$
It is clear that the assumptions on $S_{1}$ and $S_{2}$ cannot be omitted.
Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ and $n\in\mathbb{N}_{0}$. We obtain a sufficient condition for the lightness of the weight $h$ in terms of its values. We denote by $V(h,n)$ the set of vertices in $K_{\mathbb{N}}$ reachable from $1$ by a walk of length at most $n$ and with nonzero weight.
@@
Let $X\subset W(n)$ be finite. We may assume that walks in $X$ have only edges with nonzero weights. We decompose every walk $w\in X$ as $w=w^{\prime}\,\ell(w)$, where $w^{\prime}\in W(n-1)$ and $\ell(w)$ is the last vertex. Let $Y$ be the set of walks $w^{\prime}$. For $w^{\prime}\in Y$ let
@@
We get the bound
@@
Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$. We define $V(h)=\bigcup_{n\geq 0}V(h,n)$. It is the set of vertices in $K_{\mathbb{N}}$ reachable from $1$ by walks with nonzero weights. We define the graph
@@
as the connected component of $1$ in the subgraph of $K_{\mathbb{N}}$ formed by edges with nonzero weight. It follows that if $G(h)$ has finite degrees then $h$ is slim and hence light.
@@
Proof. There is a walk $w_{0}=\langle u_{0},\dots,u_{n}\rangle$ in $G(h)$ such that $u_{0}=1$ and $u_{n}=u$. Let $c=h(w_{0})$ ($\neq 0$) and $W$ be the set of walks $\langle u_{0},\dots,u_{n},u_{n+1}\rangle$ in $K_{\mathbb{N}}$. Thus vertices $u_{0}$, $\dots$, $u_{n}$ are fixed and $u_{n+1}$ runs in $\mathbb{N}\setminus\{u_{n}\}=\mathbb{N}\setminus\{u\}$. The series $\sum_{w\in W}h(w)$ is a subseries of $\sum_{w\in W(n+1)}h(w)$ and therefore absolutely converges. By Proposition 2.2,
@@
We survey some results on generating functions. We work with univariate formal power series
@@
Neutral elements are $0=0+0x+0x^{2}+\dots$ and $1=1+0x+0x^{2}+\cdots$. Units in $\mathbb{C}[[x]]$ are exactly the formal power series $\sum_{n\geq 0}a_{n}x^{n}$ with $a_{0}\neq 0$.
@@
If $U(1)_{2}$ exists, then $U(1)_{1}$ exists and $U(1)_{1}=U(1)_{2}$, but the opposite implication is, in general, not true. For example, if
@@
then $U(1)_{1}$ exists and $U(1)_{1}=\log 2$, but $U(1)_{2}$ does not exist. But if
@@
If $U(x)=\sum_{n\geq 0}u_{n}x^{n}\in\mathbb{C}[[x]]$ and $U(1)_{1}$ exists, then the power series $\sum_{n\geq 0}u_{n}x^{n}$ absolutely converges for every $x\in\mathbb{C}$ with $|x|<1$.
In fact, for the absolute convergence it suffices (and is necessary) by Hadamard’s formula that $\limsup_{n\to\infty}|u_{n}|^{1/n}\leq 1$.
@@
Quantities $U(1)_{1}$ are handled by means of the classical theorem due to N. Abel, see [1] and [28, Chapter II.7]. We state it for the radius of convergence $R\geq 1$.
@@
Let $U(x),V_{1}(x),\dots,V_{n}(x)\in\mathbb{C}[[x]]$, $n\in\mathbb{N}$. If $U(x)=\prod_{i=1}^{n}V_{i}(x)$ and $U(1)_{1}$, $V_{1}(1)_{1}$, $\dots$, $V_{n}(1)_{1}$ exist, then
@@
Proof. By Propositions 2.4, 2.6 and 2.13, and by the assumptions,
$F_{U}(x)=\prod_{i=1}^{n}F_{V_{i}}(x)$ for every $x\in\mathbb{C}$ with $|x|<1$
(see proofs of Propositions 2.25 and 2.27 for details of similar arguments). Theorem 2.16 and properties of limits of functions give
@@
Proof. In $\mathbb{C}[[x]]$ we have the identity $U(x)V(x)=1$. By Propositions 2.4 and 2.5, $F_{U}(x)F_{V}(x)=1$ for $x\in[0,1)$. Thus $F_{V}(x)\neq 0$ and $F_{U}(x)=\frac{1}{F_{V}(x)}$ for every $x\in[0,\,1)$. Two applications of Theorem 2.17 and properties of limits of functions give
@@
Proof. In $\mathbb{C}[[x]]$ we have the identity $U(x)^{2}\cdot W(x)=W(x)-V(x)$. Let T(x)=W(x)V(x)T(x)=W(x)-V(x). By Proposition 2.18, $T(1)_{1}$ exists and $T(1)_{1}=W(1)_{1}-V(1)_{1}$. By Proposition 2.19 we have $(U(1)_{1})^{2}\cdot W(1)_{1}=T(1)_{1}=W(1)_{1}-V(1)_{1}$. Division by $W(1)_{1}$ yields the stated formula. $\Box$
@@
$(U(1)_{1})^{2}$ $=$ ${\textstyle\big(\lim_{x\to 1}F_{U}(x)\big)^{2}=\lim_{x\to 1}F_{U}(x)^{2}=\lim_{x\to 1}\big(1-\frac{F_{V}(x)}{F_{W}(x)}\big)}$
$=$ ${\textstyle 1-\frac{\lim_{x\to 1}F_{V}(x)}{\lim_{x\to 1}F_{W}(x)}=1-\frac{a}{\lim_{n\to\infty}\sum_{j=0}^{n}w_{j}}=1-\frac{a}{+\infty}=1}\,.$
@@
We obtain analogs of the previous results in absolute convergence.
@@
By Proposition 2.5 it absolutely converges and has sum $V(1)_{2}W(1)_{2}$. On the other hand, $S$ has the grouping $\sum_{n\geq 0}\sum_{j=0}^{n}v_{j}w_{n-j}$ (the inner $\sum$ means sum and the outer $\sum$ means series) and $\sum_{j=0}^{n}v_{j}w_{n-j}=u_{n}$. By Proposition 2.4, $S$ has the sum $u_{0}+u_{1}+\dots=U(1)_{2}$. Thus $U(1)_{2}=V(1)_{2}W(1)_{2}$.